# DISCO: Diversifying Sample Condensation for Efficient Model Evaluation

### v1

@@ Abstract
Evaluating modern machine learning models has become prohibitively expensive. Benchmarks such as LMMs-Eval and HELM demand thousands of GPU hours per model. Costly evaluation reduces inclusivity, slows the cycle of innovation, and worsens environmental impact. The typical approach follows two steps. First, select an anchor subset of data. Second, train a mapping from the accuracy on this subset to the final test result. The drawback is that anchor selection depends on clustering, which can be complex and sensitive to design choices. We argue that promoting diversity among samples is not essential; what matters is to select samples that maximise diversity in model responses. Our method, Diversifying Sample Condensation (DISCO), selects the top-k samples with the greatest model disagreements. This uses greedy, sample-wise statistics rather than global clustering. The approach is conceptually simpler. From a theoretical view, inter-model disagreement provides an information-theoretically optimal rule for such greedy selection. DISCO shows empirical gains over prior methods, achieving state-of-the-art results in performance prediction across MMLU, Hellaswag, Winogrande, and ARC.
@@ 1 Introduction
We validate these ideas through Diversifying Sample Condensation (DISCO). DISCO selects a small, informative subset of evaluation samples by focusing on model disagreement. Disagreement is measured by predictive diversity scoring (PDS, Rubinstein et al. (2024)), originally proposed for out-of-distribution detection. A simple metamodel then predicts benchmark performance directly from the model signatures on this subset. We evaluate DISCO in both language and vision domains. On MMLU, for example, DISCO reduces evaluation cost by 99.3% with only 1.07 percentage points of error. Compared with prior methods such as Anchor Points (Vivek et al., 2023), TinyBenchmarks (Polo et al., 2024), and Metabench (Kipnis et al., 2024), DISCO achieves a stronger efficiency–precision trade-off.
@@ 2 Related work
Label-efficient evaluation. In the pre-LLM context, labelling a test set used to be a cost bottleneck for evaluation. In this context, the concept of “active testing” has been explored, where labelling budget is maximally assigned to information-rich samples (Majumdar & Niksic, 2017; Ji et al., 2021; Deng & Zheng, 2021; Kossen et al., 2021; Hu et al., 2023; Kossen et al., 2022; Huang et al., 2024). In our case, we are concerned with the inference costs of evaluation. As such, active testing approaches are not directly applicable, as they require a full inference over the test set to identify informative samples to label.
Efficient benchmarking. In the LLM era, benchmarks have diversified to measure multiple capabilities and styles of model behaviours. Researchers have proposed strategies to build an efficient benchmark in the first place (Prabhu et al., 2024; Perlitz et al., 2023; Rädsch et al., 2025). There were attempts to compress multiple benchmarks, measuring an array of capabilities of LLMs, into a single one by eliminating redundancies (Zhao et al., 2024; Yuan et al., 2025). Others have focused on selection of small, informative subsets, also known as “Anchor point” approaches (Vivek et al., 2023; Polo et al., 2024; Li et al., 2025; Gupta et al., 2025; Kipnis et al., 2024). Given an entire dataset, they compute a small subset of data points according to the representativeness criterion, determined through the correctness patterns of a large number of source models. Afterwards, the target model performance is estimated based on weighted accuracy computation on the selected subset. In particular, tinyBenchmarks (Polo et al., 2024) and Metabench (Kipnis et al., 2024) have adopted Item Response Theory (IRT) (Lord & Novick, 2008) to estimate model performance in a principled manner. Hofmann et al. (2025) proposed an IRT-based approach to LLM evaluation that selects anchor points dynamically for each model, guided by its predictions on previously chosen anchors. To address the growing number of methods for efficient LLM evaluation, Zhang et al. (2025) recently introduced a large-scale benchmark. In this work, we suggest reconsidering the approaches for both anchor point selection and target model performance estimation, as explained below.
@@ 3 Problem
Our task is the estimation of model performance on a benchmark. Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a predictive model over a dataset $\mathcal{D}:=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$ sampled iid from some distribution. We are interested in estimating the model performance on the dataset $r(f,\mathcal{D})$. An example metric for model performance is accuracy: for a probabilistic classifier $f:\mathcal{X}\rightarrow[0,1]^{C}$, accuracy is defined as $\frac{1}{N}\sum_{i}\mathbf{1}\left(\arg\max_{c}f_{c}(x_{i})=y_{i}\right)$.
We are interested in estimating ${r}(f,\mathcal{D})$ in a cost-effective way. We seek ways to sample a subset of size $K\ll N$ from the original set $\mathcal{D}$ to estimate ${r}(f,\mathcal{D})$. The overall problem is described in Figure 2. An integral ingredient for both prior works and ours is the set of source models $\mathcal{F}=\{f^{1},\ldots,f^{M}\}$, a held-out set of models whose ground-truth performances are known. We define the target models as the models whose performances we aim to estimate.
@@ 4.1.1 Prior selection methods
Best for validation Kipnis et al. (2024) finds an anchor set $A$ through an iterative search. The algorithm first generates a large number of candidate anchor sets, $\{A_{1},\dots,A_{P}\}$, by uniformly sampling from the full dataset $\mathcal{D}$. For each candidate set $A_{p}$, a simple scalar-to-scalar regression model, $g_{p}$, is trained on the source models $\mathcal{F}$. This model learns to map the performance on the subset, $r(f,A_{p})$, to the known ground-truth performance on the full dataset, $r(f,\mathcal{D})$. Each trained regressor $g_{p}$ is subsequently evaluated on a held-out validation set of models. The final anchor set $A$ is selected as the candidate $A_{p}$ whose corresponding regressor $g_{p}$ yields the lowest prediction error (e.g., RMSE) on this validation set.
How DISCO differs. Unlike clustering, we use sample-wise statistic to determine samples with maximal information content. This greatly simplifies the sampling procedure. We exploit the model diversity, not model confidence or correctness. A set of models can be highly confident and diverse at the same time. We argue that inputs that induce model diversity are more useful for performance prediction.
@@ 4.1.2 DISCO selection
We now present our selection method. We hypothesise that samples that produce the greatest diversity in model outputs carry most information about the model performance (see Figure 1). In this part, we explain how we identify such samples in the test dataset. Our sample selection strategies are illustrated in Figure 3. The main approach in Diversifying Sample Condensation (DISCO) is to select a subset $\mathcal{D}_{\text{{DISCO}}}$ of the original evaluation set $\mathcal{D}$ by sampling the top-k samples based on disagreement score, such as Predictive Diversity Score (PDS) from Equation 1. This follows the intuition shown in Figure 1.
@@
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i}^{N}$ be a test set and $m\sim\mathrm{Unif}\{1,\dots,M\}$ be the index of a uniformly chosen model. Let $f_{c}^{m}(x_{i})\in[0,1]$ be the predictive probability for class $c$ of model $f^{m}$ on input $x_{i}$. We write $\widehat{y}_{i}^{m}$ for the categorical random variable following $\operatorname{Cat}(f_{1}^{m}(x_{i}),\ldots,f_{C}^{m}(x_{i}))$. Define ensemble mean prediction to be $\widebar{f}_{c}(x_{i}):=\mathbb{E}_{m}[f^{m}_{c}(x_{i})]$ for each class $c$ and define corresponding prediction random variable as $\widehat{y}_{i}$ following $\operatorname{Cat}(\widebar{f}_{1}(x_{i}),\ldots,\widebar{f}_{C}(x_{i}))$. Let $S(m)=S(f^{m},\mathcal{D})$ denote a function of model $m$ and dataset $\mathcal{D}$, such as model accuracy, that is injective with respect to $m$. We make additional Assumptions A1A2 in appendix. Then,
@@
where ${H}(\cdot)$ is entropy, $\mathrm{MI}(\cdot)$ is mutual information, and $\operatorname{JSD}(\cdot)$ is generalised Jensen-Shannon Divergence for multiple distributions (Fuglede & Topsoe, 2004).
See proof in Appendix A.
We conclude that the sample $i$ conveying the greatest level of information for the prediction of $S(m)$ (e.g. model accuracy) is the one with greatest $\operatorname{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right)$. This generalised Jensen-Shannon divergence translates to the diversity of distributions (Fuglede & Topsoe, 2004). Based on the insight that model diversity matters for performance prediction, we also consider an alternative measure that measures the model diversity: predictive diversity score (PDS) (Rubinstein et al., 2024). It is more interpretable, as it is a continuous generalisation of the number of unique argmax category predictions among $M$ source models:
@@ 4.2.1 Prior prediction methods
ability-IRT Kipnis et al. (2024) is a two-stage method that uses the anchor set $A$ as a diagnostic tool rather than just a miniature test. First, it uses a pre-calibrated IRT model to estimate a latent “ability” score, $\hat{\theta}^{f}$, from the target model’s pattern of correct and incorrect responses on $A$. Second, a pre-trained regressor, $g$, predicts the final performance $S^{f}$ using both the simple anchor set accuracy $\hat{S}^{f}_{A}$ and this more informative ability score $\hat{\theta}^{f}$ as input features. The final prediction is given by $S^{f}=g(\hat{S}^{f}_{A},\hat{\theta}^{f})$, leveraging a deeper measure of the model’s capability to improve the estimate.
@@ 4.2.2 DISCO prediction
Given a smaller set of test dataset $\mathcal{D}_{\text{{DISCO}}}$, we estimate the performance of a model $f$ as closely as possible to the true full test performance ${r}(f,\mathcal{D})$. We deliberately opt for simple approaches here, in order to make a point that simple is best; we also compare against a rather complex prior work and show that our simple method wins against the stolen. Our performance prediction framework is depicted in Figure 3.
@@
Such function signature may have a large dimensionality, as it is the product of model output dimensionality (e.g. 1000 for ImageNet) and the number of selected samples $|\mathcal{D_{\text{DISCO}}}|$ (e.g. can go up to 50k for ImageNet validation set). To reduce the storage burden and improve generalizability, we consider applying a dimenionality reduction technique based on principal component analysis (PCA): $Q\circ{f}(\mathcal{D}_{\text{{DISCO}}})$.
@@
Parametric mapping. We also consider a parametric prediction variant. A single parametric mapping $R$ is trained for the prediction of model performance. As the training set, we use $M$ model signatures $Q\circ{f}_{1}(\mathcal{D}_{\text{{DISCO}}}),\ldots Q\circ{f}_{M}(\mathcal{D}_{\text{{DISCO}}})$ for $\mathcal{F}$ as the training set for the regression problem of training a mapping $R(\cdot)$ to let $R\circ Q\circ{f}_{m}(\mathcal{D}_{\text{{DISCO}}})$ approximate $\widehat{r}(f,\mathcal{D})$. The predictor $R$ can be implemented using a neural network, linear regression, or a Random Forest, for example.
@@ 5 Experiments
In this section, we introduce the experimental setup, present the main results of Diversifying Sample Condensation (DISCO) in language domain (§5.1), analyse contributing factors (§5.2), and demonstrate that the method is domain-agnostic and can also be successfully applied to the vision domain (§5.3).
Datasets. We evaluate DISCO on two widely used language modeling benchmarks: MMLU (Hendrycks et al., 2021) and HellaSwag (Zellers et al., 2019). MMLU covers 57 tasks about world knowledge and problem-solving ability, and HellaSwag focuses on commonsense natural language inference.
Models. Building on the TinyBenchmarks framework (Polo et al., 2024), we evaluate 424 large language models (LLMs) from Hugging Face’s Open LLM Leaderboard (Fourrier et al., 2024). The models cover GPT- (Radford et al., 2019), LLaMA- (Touvron et al., 2023), DeepSeek- (DeepSeek-AI et al., 2025), and BERT-style (Devlin et al., 2019) architectures, with model sizes ranging from 1.3 billion to 32 billion parameters.
Model split. DISCO is based on a meta-model approach where a predictor is constructed based on the model signatures of a pool of source models $\mathcal{F}$ and tested on a disjoint set of models. This approach has traditionally been criticised for its dependency on the set of existing models: the approach may fail to retain performance with unforeseen changes in future models. To address this concern, we introduce the chronological split, where the source models $\mathcal{F}$ consist of models published before January 13, 2024 and the meta test set consists of models after the cutoff date. The train-test ratio is 9:1.

### v2

@@ Abstract
Evaluating modern machine learning models has become prohibitively expensive. Benchmarks such as LMMs-Eval and HELM demand thousands of GPU hours per model. Costly evaluation reduces inclusivity, slows the cycle of innovation, and worsens environmental impact. To address the growing cost of standard evaluation, new methods focused on efficient evaluation have started to appear. The typical approach follows two steps. First, select an anchor subset of data. Second, train a mapping from the accuracy on this subset to the final test result. The drawback is that anchor selection depends on clustering, which can be complex and sensitive to design choices. We argue that promoting diversity among samples is not essential; what matters is to select samples that maximise diversity in model responses. Our method, Diversifying Sample Condensation (DISCO), selects the top-k samples with the greatest model disagreements. This uses greedy, sample-wise statistics rather than global clustering. The approach is conceptually simpler. From a theoretical view, inter-model disagreement provides an information-theoretically optimal rule for such greedy selection. DISCO shows empirical gains over prior methods, achieving state-of-the-art results in performance prediction across MMLU, Hellaswag, Winogrande, and ARC.
@@ 1 Introduction
We validate these ideas through Diversifying Sample Condensation (DISCO). DISCO selects a small, informative subset of evaluation samples by focusing on model disagreement. Disagreement is measured by predictive diversity scoring (PDS, Rubinstein et al. (2024)), originally proposed for out-of-distribution detection. A simple metamodel then predicts benchmark performance directly from the model signatures on this subset. We evaluate DISCO in both language and vision domains. On MMLU, for example, DISCO reduces evaluation cost by 99.3% (see Appendix B.4) with only 1.07 percentage points of error. Compared with prior methods such as Anchor Points (Vivek et al., 2023), TinyBenchmarks (Polo et al., 2024), and Metabench (Kipnis et al., 2024), DISCO achieves a stronger efficiency–precision trade-off.
@@ 2 Related work
Label-efficient evaluation. In the pre-LLM context, labelling a test set used to be a cost bottleneck for evaluation. In this context, the concept of “active testing” has been explored, where labelling budget is maximally assigned to information-rich samples (Majumdar and Niksic, 2017; Ji et al., 2021; Deng and Zheng, 2021; Kossen et al., 2021; Hu et al., 2023; Kossen et al., 2022; Huang et al., 2024; Fogliato et al., 2024). In our case, we are concerned with the inference costs of evaluation. As such, active testing approaches are not directly applicable, as they require a full inference over the test set to identify informative samples to label.
Efficient benchmarking. In the LLM era, benchmarks have diversified to measure multiple capabilities and styles of model behaviours. Researchers have proposed strategies to build an efficient benchmark in the first place (Perlitz et al., 2023; Rädsch et al., 2025). There were attempts to compress multiple benchmarks, measuring an array of capabilities of LLMs, into a single one by eliminating redundancies (Kipnis et al., 2024; Zhao et al., 2024; Yuan et al., 2025). Others have focused on selection of small, informative subsets, also known as “Anchor point” approaches (Vivek et al., 2023; Polo et al., 2024; Li et al., 2025; Gupta et al., 2025). Given an entire dataset, they compute a small subset of data points according to the representativeness criterion, determined through the correctness patterns of a large number of source models. Subsequently, target model performance is estimated based on weighted accuracy computed on the selected subset. In particular, tinyBenchmarks (Polo et al., 2024) have adopted Item Response Theory (IRT) (Lord and Novick, 2008) to estimate model performance in a principled manner. Hofmann et al. (2025) proposed an IRT-based approach to LLM evaluation that selects anchor points dynamically for each model, guided by its predictions on previously chosen anchors. To address the growing number of methods for efficient LLM evaluation, Zhang et al. (2025) recently introduced a large-scale benchmark. In this work, we adopt approaches from the black-box model analysis techniques, explained below.
@@ 3 Problem
Our task is the estimation of model performance on a benchmark. Let $f:\mathcal{X}\rightarrow\mathcal{Y}$ be a predictive model over a dataset $\mathcal{D}:=\{(x_{1},y_{1}),\ldots,(x_{N},y_{N})\}$ sampled iid from some distribution. We are interested in estimating the model performance on the dataset $S^{f}_{\mathcal{D}}$. An example metric for model performance is accuracy: for a probabilistic classifier $f:\mathcal{X}\rightarrow[0,1]^{C}$, accuracy is defined as $\frac{1}{N}\sum_{i}\mathbf{1}\{\arg\max_{c}f_{c}(x_{i})=y_{i}\}$.
[…]
@@ 4.1.1 Prior selection methods
Best for validation Kipnis et al. (2024) finds an anchor set $A$ through an iterative search. The algorithm first generates a large number of candidate anchor sets, $\{A_{1},\dots,A_{P}\}$, by uniformly sampling from the full dataset $\mathcal{D}$. For each candidate set $A_{p}$, a simple scalar-to-scalar regression model, $g_{p}$, is trained on the source models $\mathcal{F}$. This model learns to map the performance on the subset, $S^{f}_{A_{p}}$, to the known ground-truth performance on the full dataset, $S^{f}_{\mathcal{D}}$. Each trained regressor $g_{p}$ is subsequently evaluated on a held-out validation set of models. The final anchor set $A$ is selected as the candidate $A_{p}$ whose corresponding regressor $g_{p}$ yields the lowest prediction error (e.g., RMSE) on this validation set.
How DISCO differs. Unlike clustering, we use sample-wise statistics to determine samples with maximal information content. This greatly simplifies the sampling procedure. We exploit the model diversity, not model confidence or correctness. A set of models can be highly confident and diverse at the same time. We argue that inputs that induce model diversity are more useful for performance prediction.
@@ 4.1.2 DISCO selection
We now present our selection method. In this part, we explain how we identify such samples in the test dataset. Our sample selection strategies are illustrated in Figure 3. The main approach in Diversifying Sample Condensation (DISCO) is to select a subset $\mathcal{D}_{\text{{DISCO}}}$ of the original evaluation set $\mathcal{D}$ by sampling the top-k samples based on disagreement score, such as PDS. This follows the intuition shown in Figure 1.
@@
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i}^{N}$ be a test set and $m\sim\mathrm{Unif}\{1,\dots,M\}$ (A1) be the index of a uniformly chosen model. Let $f_{c}^{m}(x_{i})\in[0,1]$ be the predictive probability for class $c$ of model $f^{m}$ on input $x_{i}$. We write $\widehat{y}_{i}^{m}$ for the categorical random variable following $\operatorname{Cat}(f_{1}^{m}(x_{i}),\ldots,f_{C}^{m}(x_{i}))$. Define ensemble mean prediction to be $\widebar{f}_{c}(x_{i}):=\mathbb{E}_{m}[f^{m}_{c}(x_{i})]$ for each class $c$ and define corresponding prediction random variable as $\widehat{y}_{i}$ following $\operatorname{Cat}(\widebar{f}_{1}(x_{i}),\ldots,\widebar{f}_{C}(x_{i}))$. Let $S(m)=S(f^{m},\mathcal{D})$ denote a function of model $m$ and dataset $\mathcal{D}$, such as model accuracy, that is injective with respect to $m$. Assume that the only randomness in $\widehat{y}^{m}$ comes from $m$ (A2). Then,
@@
where ${H}(\cdot)$ is entropy, $\mathrm{MI}(\cdot)$ is mutual information, and $\operatorname{JSD}(\cdot)$ is generalised Jensen-Shannon Divergence for multiple distributions (Fuglede and Topsoe, 2004).
See proof in Appendix G.
We conclude that the sample $i$ conveying the greatest level of information for the prediction of $S(m)$ (e.g., model accuracy) is the one with the greatest $\operatorname{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right)$. This generalised Jensen-Shannon divergence translates to the diversity of distributions (Fuglede and Topsoe, 2004). Based on the insight that model diversity matters for performance prediction, we also consider an alternative measure that measures the model diversity: predictive diversity score (PDS) (Rubinstein et al., 2024). It is more interpretable, as it is a continuous generalisation of the number of unique argmax category predictions among $M$ source models:
@@ 4.2.1 Prior prediction methods
ability-IRT Kipnis et al. (2024) is a two-stage method that uses the anchor set $A$ as a diagnostic tool rather than just a miniature test. First, it uses a pre-calibrated IRT model to estimate a latent “ability” score, $\hat{\theta}^{f}$, from the target model’s pattern of correct and incorrect responses on $A$. Second, a pre-trained regressor, $g$, predicts the final performance $S_{\mathcal{D}}^{f}$ using both the simple anchor set accuracy $\hat{S}^{f}_{A}$ and this more informative ability score $\hat{\theta}^{f}$ as input features. The final prediction is given by $S_{\mathcal{D}}^{f}=g(\hat{S}^{f}_{A},\hat{\theta}^{f})$, leveraging a deeper measure of the model’s capability to improve the estimate.
@@ 4.2.2 DISCO prediction
Given a smaller set of test dataset $\mathcal{D}_{\text{{DISCO}}}$, we estimate the performance of a model $f$ as closely as possible to the true full test performance $S^{f}_{\mathcal{D}}$. We deliberately opt for simple approaches here, in order to make a point that simple is best; we also compare against a rather complex prior work and show that our simple method outperforms it. Our performance prediction framework is depicted in Figure 3.
@@
Such a function signature may have high dimensionality, as it is the product of model output dimensionality (e.g., 1000 for ImageNet) and the number of selected samples $|\mathcal{D_{\text{DISCO}}}|$ (e.g., can go up to 50k for ImageNet validation set). To reduce the storage burden and improve generalizability, we consider applying a dimensionality reduction technique based on principal component analysis (PCA): $Q\circ{f}(\mathcal{D}_{\text{{DISCO}}})$.
@@
Parametric mapping. We also consider a parametric prediction variant. A single parametric mapping $R$ is trained for the prediction of model performance. As the training set, we use $M$ model signatures $Q\circ{f}_{1}(\mathcal{D}_{\text{{DISCO}}}),\ldots Q\circ{f}_{M}(\mathcal{D}_{\text{{DISCO}}})$ for $\mathcal{F}$ as the training set for the regression problem of training a mapping $R(\cdot)$ to let $R\circ Q\circ{f}_{m}(\mathcal{D}_{\text{{DISCO}}})$ approximate $\widehat{S}^{f}_{\mathcal{D}}$. The predictor $R$ can be implemented using a neural network, linear regression, or a Random Forest, for example.
@@ 5 Experiments
In this section, we introduce the evaluation protocol (§5.1) and the experimental setup (§5.2), present the main results of Diversifying Sample Condensation (DISCO) in language domain (§5.3), analyse contributing factors (§5.4), and demonstrate that the method is domain-agnostic and can also be successfully applied to the vision domain (§5.5).
5.1 Evaluation protocol
To ensure a fair comparison, all methods follow an identical evaluation protocol, i.e., they use the same ingredients and perform the same sequence of steps during the training and testing stages.
Training. Following § 4, select anchor datapoints and train the performance predictor.
Input: source models $\mathcal{F}=\{f^{1},\ldots,f^{M}\}$, full test dataset $\mathcal{D}=(\mathcal{D}_{x},\mathcal{D}_{y})$ with questions $\mathcal{D}_{x}$ and ground truth answers $\mathcal{D}_{y}$, parameter K{K}.
Output: set of anchor datapoints $A_{K}$, predictor $R$.
  1. 1.
    Evaluate source models $\mathcal{F}$ on $\mathcal{D}_{x}$ and obtain model outputs
    $\mathcal{F}(\mathcal{D}_{x})=\{f(x):x\in\mathcal{D}_{x},f\in\mathcal{F}\}$.
  2. 2.
    Calculate their full test performance (e.g., accuracy) $S^{\mathcal{F}}_{\mathcal{D}}=\{S^{f}_{\mathcal{D}}:f\in\mathcal{F}\}$ based on $\mathcal{F}(\mathcal{D}_{x})$ and $\mathcal{D}_{y}$.
  3. 3.
    Use $\mathcal{F}(\mathcal{D}_{x})$ and optionally $\mathcal{D}_{y}$ to select a set $A_{K}\subseteq\mathcal{D}_{x}$ of $K$ anchor datapoints with respective selection method (e.g., PDS/JSD, IRT, etc.) explained in § 4.1.
  4. 4.
    Train a predictor $R$ (e.g., Random Forest, gp-IRT, ability-IRT, etc.) explained in § 4.2 to predict full test performance from model’s outputs on anchor datapoints
    $f(A_{K})=\{f(x):x\in A_{K}\}$, such that ${S}^{f}_{D}\approx\widehat{S}^{f}_{D}=R(f(A_{K})),\forall f\in\mathcal{F}$.
Testing. Test the performance predictor that is trained as explained above.
[…]
Output: performance of the efficient evaluation method.
  1. 1.
    Evaluate target models $\mathcal$$\tilde${F}} on anchor points $A_{K}$} and obtain their outputs
[…]
  2. 2.
    Use predictor $R$ to estimate the ground truth performances of the target models
[…]
  3. 3.
    Calculate performance (e.g., MAE, Spearman rank correlation, etc.) of the efficient evaluation method by comparing $S^{\tilde{F}}_{\mathcal{D}}$ and $\widehat{S}^{\tilde{F}}_{\mathcal{D}}$.
5.2 Setup
We describe our experimental setup: the datasets, metrics, models, and model splits.
Datasets. We evaluate DISCO on four widely used language modeling benchmarks: MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), and ARC (Clark et al., 2018). Details on the benchmarks can be seen in Appendix A.